OpenStudy (anonymous):

How does changing the function from f(x) = 2 sin 4x to g(x) = 2 sin 4x + 3 affect the range of the function? The function shifts up 3 units, so the range changes from −2 to 2 in f(x) to 1 to 5 in g(x). The function shifts up 3 units, so the range changes from −1 to 1 in f(x) to 2 to 4 in g(x). The function shifts up 4 units, so the range changes from −2 to 2 in f(x) to 2 to 6 in g(x). The function shifts up 4 units, so the range changes from −1 to 1 in f(x) to 3 to 5 in g(x).

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@Destinymasha @nincompoop

OpenStudy (mathmale):

Have you tried graphing both functions? If you have a graphing calculator, that should be relatively easy to do. Once you have the pertinent graphs, you can easily compare them visually. Find the range of the original function. Then, modify this range by adding 3 to both upper and lower limit on y.

OpenStudy (anonymous):

What are the period and phase shift for f(x) = 5 tan(2x − π)? period: π; phase shift: x = pi over two period: π; phase shift: x = negative pi over two period: pi over two; phase shift: x = negative pi over two period: pi over two; phase shift: x = pi over two

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Have you found the answer for the first question?

OpenStudy (anonymous):

Yeah! @ranga

OpenStudy (ranga):

f(x) = 5 tan(2x − π) Put it in form: A*tan(B(x - C)) where A is amplitude; period = pi/B; phase shift = C. f(x) = 5 tan(2x − π) The coefficient of x is 2. You have to factor out the 2 to make the coefficient of x one. f(x) = 5 tan(2(x − π/2)). A = 5, B = 2, C = π/2 Period = pi/B = ? Phase shift = C = ?

OpenStudy (anonymous):

What is the rate of change for f(x) = 4 sin x − 2 on the interval from x = 0 to x = pi over 2? pi over 4 4 over pi pi over 8 8 over pi

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Average rate of change of f(x) in the interval [a,b] = (f(b) - f(a)) / (b - a)

OpenStudy (anonymous):

Would it be set up as f(pi/2) - f(0)/(pi/2)

OpenStudy (ranga):

Yes. \[ \Large \frac{f(\pi / 2) - f(0)}{\pi /2} = \frac{\{4\sin(\pi / 2) - 2\}- \{4\sin(0) - 2\}}{\pi /2} = \ ? \]